# Stabilizing Thompson Sampling with Point Null Bayesian Response-Adaptive Randomization

### v1

@@ 7 Discussion
In this paper we have proposed a modification of standard Bayesian RAR (Thompson sampling) via recasting of the problem in a Bayesian hypothesis testing framework and the introduction of a point null hypothesis. While the plausibility of point null hypotheses is often a matter of philosophical debate (see e.g., Berger and Delampady,, 1987; Ly and Wagenmakers,, 2022), this method is useful in the RAR setting, as it can interpolate between equal randomization and Thompson sampling by changing the prior probability of the null hypothesis $\Pr(H_{0})$. This allows experimenters to balance patient benefit with classical operating characteristics, such as power, type I error rate, bias, and coverage. For large values of $\Pr(H_{0})$, we observed behaviors and operating characteristics similar to those obtained with ad hoc modifications of Thompson sampling, such as capping, burn-ins, and power transformations. Our method is implemented in the free and open source R package brar for binomial outcomes and for data summarized by approximately normal effect estimates. The latter makes the method applicable to many settings, for example, settings where treatment effects are estimated with regression analyses.
@@
One advantage of our framework is that the randomization probabilities coherently correspond with the available statistical evidence (in the form of Bayes factors) and beliefs (in the form of posterior probabilities). In principle, both could also be used as decision-making tools instead of relying on frequentist test criteria. For instance, if the posterior probability of a treatment’s superiority is greater than 0.99, say, a study could be stopped. This also makes sense from the perspective that it is unnatural to randomize participants with extreme randomization probabilities that are associated with such high posterior probabilities.
Although we conducted a simulation study to understand the method’s basic behavior, more realistic and comprehensive evaluations are needed to understand its applicability in real-world conditions. For example, the method needs to be evaluated in combination with futility stopping (e.g., dropping of treatment arms which are shown to be ineffective at an interim analysis). Another issue to consider is how to select the prior probability of the null hypothesis. In our simulation study, we found that setting a value of $\Pr(H_{0})=0.75$ mitigated many of the issues with Thompson sampling. However, other choices could be considered, such as setting a higher value. Similar considerations apply to the prior distributions of the parameters, as we did not assess the effect of varying these on the operating characteristics. For instance, rather than setting the correlation of the multivariate normal prior to achieve equal randomization probabilities, it could be beneficial to set it so that the prior probability of the control being superior is always $\Pr(H_{-})=0.5$ and the remaining probability is distributed equally among the treatments. Future work may therefore investigate whether a more efficient RAR procedure can be obtained by specifying a certain prior distribution. Finally, in many clinical trial settings, RAR methods are not directly applicable because outcomes such as death may only be observed after long follow-up periods, by which time recruitment and randomized allocation will already have finished. An alternative could be to perform point null Bayesian RAR with an informative surrogate outcome that is sooner observed than the primary outcome (Gao et al.,, 2024).
Acknowledgments
We thank František Bartoš and Małgorzata Roos for valuable comments on drafts of the manuscript. The acknowledgment of these individuals does not imply their endorsement of the paper.
Conflict of interest
We declare no conflict of interest.
Software and data
Code and data to reproduce our analyses are openly available at https://github.com/SamCH93/brar. A snapshot of the repository at the time of writing is available at https://doi.org/10.5281/zenodo.17248628. We used the statistical programming language R version 4.5.0 (2025-04-11) for analyses (R Core Team,, 2024) along with the ggplot2 (Wickham,, 2016), dplyr (Wickham et al.,, 2023), SimDesign (Chalmers and Adkins,, 2020), mvtnorm (Genz and Bretz,, 2009), ggh4x (van den Brand,, 2024), ggpubr (Kassambara,, 2023), and knitr (Xie,, 2015) packages.
@@ Appendix A The R package brar
Our R package can be installed by running remotes::install_github(repo = "SamCH93/brar", subdir = "package") in an R session (requires the remotes package, which is available on CRAN). The main functions of the package are brar_normal and brar_binomial, which implement the approximate normal method from Section 3 and the exact binomial method from Section 4. The following code chunk illustrates how the latter function can be used.
@@
Appendix B Simulation study
@@
B.1 Aims
The aim of the simulation study is to evaluate the design characteristics of the newly proposed point null Bayesian RAR approach, and compare it to existing methods.
B.2 Data-generating mechanism
@@
Parameters were chosen similar to the simulation study from Robertson et al., (2023). We vary the sample size $n\in\{200,654\}$ to represent low and high powered studies, the number of treatment groups $K\in\{1,2,3\}$, and probability in the first treatment group $\theta_{1}\in\{0.25,0.35,0.45\}$. The probability in the control group and the remaining groups is always fixed at $\theta_{C}=0.25$ and $\theta_{2}=\theta_{3}=0.3$, respectively. All these parameters are varied fully-factorially, leading to $2\times 3\times 3=18$ parameter conditions.
@@
B.3 Estimands and other targets
@@
The primary interest of this simulation study lies in assessing the patient benefit characteristics of different RAR methods. Additionally, the estimand of interest is the rate difference $\text{RD}_{1}=\theta_{1}-\theta_{C}$ and the target of interest is the null hypothesis of $\text{RD}_{1}=0$.
B.4 Methods
@@
We consider the above described point null RAR methods. The prior probability of $H_{0}$ is a tuning parameter and controls the variability of the randomization probabilities. Setting $\Pr(H_{0})=1$ produces equal randomization, whereas $\Pr(H_{0})=0$ produces Thompson sampling. We consider values of $\Pr(H_{0})\in\{0,0.25,0.5,0.75,1\}$, as well as the normal approximation and exact binomial version of RAR. For approximate normal RAR, a normal prior with mean 0, variance 1, and in case of $K>1$ a correlation of 0.5, is considered. Independent uniform priors are assigned for binomial RAR. Log odds ratios along with their covariance are estimated with logistic regression and then used as inputs for the normal RAR method, while the exact method uses success counts and sample sizes only. In case a method fails to converge, equal randomization is applied as a back-up strategy, as this mimics what an experimenter might do in practice when a RAR method fails to converge (Pawel et al.,, 2025).
@@
We also consider three modifications of these methods: In some conditions, a “burn-in” phase is carried out during which the first 50 patients are always randomized with equal probability $1/(K+1)$ to each group. For Thompson sampling ($\Pr(H_{0})=0$), we additionally consider conditions with power transformations of randomization probabilities, i.e., if $\pi_{k}$ is the randomization probability of group $k$, we take $\pi_{k}^{*}=\pi_{k}^{c}/\sum_{j\in\{C,1,\dots,K\}}\pi_{j}^{c}$. We consider $c=1/2$ and $c=i/(2n)$ with $i$ the current and $n$ the maximal sample size, which are two popular choices of the tuning parameter $c$ (Wathen and Thall,, 2017). Additionally, in some conditions “capping” is applied to Thompson sampling. That is, randomization probabilities outside the $10\%,90\%]$ interval are set to either 10% or 90%. After capping has been performed, randomization probabilities are re-normalized to sum to one (Wathen and Thall,, 2017; Lee and Lee,, 2021). This re-normalization is only performed for randomization probabilities greater than 10%, as these would otherwise be reduced again to probabilities less than 10%. In case, a re-normalized probability becomes less than 10%, it is also capped at 10% and second re-normalization performed. Finally, for equal randomization ($\Pr(H_{0})=1$), no burn-in, capping, or power transformation conditions are simulated as these manipulations have no effect.
B.5 Performance measures
Different performance measures were used. Patient benefit was quantified with:
@@ B.5 Performance measures
    where $n_{1i}$ is the number of allocations to treatment group 1 in simulation $i$. For $K=1$, this reduces to the $\hat{S}_{0.1}$ sample size imbalance measure from Robertson et al., (2023), which in turn was inspired by the performance evaluation in the simulation study from Thall et al., (2015).
@@
B.6 Computational aspects
The simulation study was run on a server running Debian GNU/Linux 13 (trixie) and R version 4.5.0 (2025-04-11). The SimDesign R package was used to organize and run the simulation study (Chalmers and Adkins,, 2020). Our newly developed brar R package was used to perform Bayesian RAR. Code and data to reproduce this simulation study are available at https://github.com/SamCH93/brar.
B.7 Results
@@
The mean rate of successes is shown in Figure LABEL:fig:successrate. It was generally the highest for Thompson sampling and lowest for equal randomization. The normal approximation and exact method produced mostly similar rates, with the normal method sometimes showing slightly higher rates (e.g., for $K=3$ and $\text{RD}_{1}=0.2$). The Thompson sampling modifications generally reduced the mean success rate, with the greatest reduction achieved by combining capping and power transformation with $c=i/(2n)$. In conditions with small sample size these rates were similar as when the prior probability was $H_{0}=0.75$, while they were lower when the sample size was larger. This makes sense as for larger sample sizes, uncapped randomization probabilities are more likely to converge to extreme ones. This is also visible in Figure LABEL:fig:extreme where more extreme randomization probabilities are observed with increasing sample size and rate difference for all methods but the ones with capping.
@@
Figure LABEL:fig:imbalance shows negative sample size imbalance as quantified by the $\hat{S}_{0.1}$ metric. Negative imbalance was the greatest for Thompson sampling and reduced when modifications were applied. Similarly, increasing the prior probability of $H_{0}$ decreased negative imbalance, in some cases even below Thompson sampling with modifications.
@@
Figures LABEL:fig:bias and LABEL:fig:coverage show empirical bias and coverage related to estimates of the rate difference between the first treatment group and the control group $\text{RD}_{1}$. In conditions where there was no difference ($\text{RD}_{1}=0$), all methods produced unbiased point estimates, though all methods but equal randomization also showed undercoverage. Bias occurred in conditions with non-zero rate differences, the bias being the greatest for Thompson sampling and decreasing to some extent when modifications were introduced or the prior probability of $H_{0}$ increased. Similarly, modifications or increasing prior probabilitities improved coverage, although they still remained suboptimal in most conditions. For large rate difference conditions, coverage was much better for the normal than the exact version of RAR. Similarly, bias was in some cases larger for the exact compared to the normal version.
@@
Figures LABEL:fig:t1e and LABEL:fig:power show empirical type I error rate and power associated with the Wald test of $\text{RD}_{1}=0$. We see that standard Thompson sampling shows an inflated type I error rate above the nominal 2.5% which is reduced to some extent by increasing either the prior probability of $H_{0}$ or applying modifications. In the same way, Thompson sampling exhibits reduced power compared to equal randomization, which is again alleviated by modifications or positive prior probability of the null hypotheses. In small sample sizes and large rate differences, power was slightly increased for the exact compared to the normal version of RAR, while for Thompson sampling the type I rate was slightly higher for the exact compared to the normal version.

### v2

@@ 7 Discussion
We have proposed a modification of Thompson sampling by recasting the problem in a Bayesian hypothesis testing framework and introducing a null hypothesis of equal effectiveness. This method can interpolate between equal randomization and Thompson sampling by changing the prior probability of the null hypothesis $\Pr(H_{0})$. This allows balancing patient benefit with inferential properties. For large values of $\Pr(H_{0})$, we observed similar characteristics as with ad hoc modifications of Thompson sampling.
@@
One advantage of the method is that randomization probabilities coherently correspond to the available statistical evidence (in the form of Bayes factors) and beliefs (in the form of posterior probabilities). Both could also serve as decision-making tools instead of frequentist criteria. For example, a study could be stopped if the posterior probability of efficacy exceeds 0.99. This also aligns with the perspective that it is unnatural to randomize patients using extreme randomization probabilities associated with such high posterior probabilities.
Although we conducted a simulation study to understand the method’s basic behavior, more evaluations are needed to understand its real-world applicability. For example, the method needs to be evaluated in combination with futility stopping (e.g., dropping of arms at interim) and under time trends (e.g., effects attenuating over time because the standard of care improves). RAR is often impractical when primary outcomes require long follow-up. An alternative could be to perform Bayesian RAR with a surrogate outcome that is sooner observed (Gao et al.,, 2024). Finally, further theoretical work is needed to understand the asymptotic behavior of the method. Most importantly, the empirically observed convergence of randomization probabilities to 50% under $H_{0}$ remains to be shown rigorously.
Regarding prior choice, $\Pr(H_{0})=0.75$ mitigated many Thompson sampling issues in our study, yet higher values may also be considered. Similar considerations apply to the prior distributions of the parameters. For instance, rather than setting the correlation of the multivariate normal prior to achieve equal randomization probabilities, it could be set so that the prior probability of the control being superior is always $\Pr(H_{0})=0.5$ and the remaining probability is equally distributed among the treatments. Similarly, concurrent allocation to the control group may be desirable (Trippa et al.,, 2012). Future work may therefore investigate whether more efficient RAR can be obtained by tuning the prior distribution.
Acknowledgments and conflict of interest
We thank František Bartoš and Małgorzata Roos for valuable comments. We thank the associate editor and two anonymous reviewers for excellent suggestions. We declare no conflict of interest.
Supplementary materials
Supplementary materials and data and code are available with this paper at the Biometrics website on Oxford Academic.
Data availability
Code and data to reproduce our results are available at https://github.com/SamCH93/brar and archived at https://doi.org/10.5281/zenodo.17248628.
@@ Appendix A The R package brar
Our R package can be installed by running install.packages("brar") in an R console. The main functions of the package are brar_normal and brar_binomial, which implement the normal method from Section 3 and the binomial method from Section 4 in the main text. The following code chunk illustrates how the latter function can be used.
@@
Appendix B Null hypothesis Bayesian RAR under multivariate normality
Suppose there are $K>1$ treatment groups and consequently $K$ effect estimates, each estimate quantifying the effect of the corresponding treatment relative to the control. A natural generalization univariate normal Bayesian RAR is to stack the estimates into a vector $\boldsymbol{\hat{\theta}}=(\hat{\theta}_{1},\dots,\hat{\theta}_{K})^{\top}$ and assume a $K$-variate normal distribution $\boldsymbol{\hat{\theta}}\mid\boldsymbol{\theta}\sim\mathrm{N}_{K}(\boldsymbol{\theta},\boldsymbol{\boldsymbol{\Sigma}})$, where $\boldsymbol{\theta}=(\theta_{1},\dots,\theta_{K})^{\top}$ is the vector of true effects and $\boldsymbol{\boldsymbol{\Sigma}}$ is the (assumed to be known) covariance matrix of $\boldsymbol{\hat{\theta}}$. For example, $\boldsymbol{\hat{\theta}}$ could be a vector of estimated regression coefficients and $\boldsymbol{\boldsymbol{\Sigma}}$ its estimated covariance matrix.
The hypotheses from Section 2 in the main text then translate into
$H_{-}\colon\max\{\theta_{1},\dots,\theta_{K}\}<0$ vs. $H_{0}\colon\theta_{1}=\dots=\theta_{K}=0$ vs. $H_{+i}\colon\theta_{i}=\max\{\theta_{1},\dots,\theta_{K}\}>0$
for $i=1,\dots,K$. All hypotheses apart from $H_{0}$ are composite and require the specification of a prior distribution. In analogy to the one treatment case, we specify a $K$-variate normal prior $\boldsymbol{\theta}\sim\mathrm{N}_{K}(\boldsymbol{\mu},\boldsymbol{\mathcal{T}})$ and truncate its support to the region of the corresponding hypothesis (e.g., for hypothesis $H_{+i}$, the space in $\mathbb{R}^{K}$ where the $i$th component is positive and larger than and all other components). As in the one group case, it seems a sensible default to center the prior on $\boldsymbol{0}$ to encode clinical equipoise. Similarly, the prior hypothesis probabilities $\Pr(H_{+i})$ for $i=1,\dots,K$ may again be specified so that the $\mathrm{N}_{K}(\boldsymbol{\mu},\boldsymbol{\mathcal{T}})$ distribution is recovered when the prior is averaged over the hypotheses. Additionally, specifying a prior covariance matrix $\boldsymbol{\mathcal{T}}$ with uniform correlation of 0.5 ensures equal treatment prior probabilities for any number of treatment groups $K$, and thus seems a sensible default in the absence of prior knowledege.
The normal-normal conjugate framework allows us to derive marginal likelihoods in closed-form. The marginal likelihood under $H_{0}$ is
while the marginal likelihood under $H_{-}$ is
with the first row encoding the constraint of $\theta_{2}$ being positive, and the second and third rows encoding the constraints of $\theta_{2}$ being larger than $\theta_{1}$ and $\theta_{3}$, respectively. As expected, for $K=1$, these marginal likelihoods reduce to the univariate ones (7) in the main text. Moreover, since they are all available in closed-form, Bayes factors, posterior probabilities, and randomization probabilities are also available in closed-form. Crucially, only the exact evaluation of multivariate normal densities and cumulative distribution functions is required, and both are efficiently implemented in statistical software (e.g., in the mvtnorm R package, Genz and Bretz,, 2009). This thus leads to an efficient Bayesian RAR method that allows experimenters to compute randomization probabilities in complex settings, for instance, multiple regression, where a full Bayesian analysis may involve additional complexities, such as priors for nuisance parameters and Markov chain Monte Carlo methods for the computation of posteriors.
Figure S1: Evolution of Bayesian RAR probabilities for 3 treatments and a control group. In each step, one observation from each group is simulated from a normal distribution with a standard deviation of 1 and corresponding mean differences as indicated in the panel titles (treatment 1 is the most effective treatment), and a mean difference effect estimate $\boldsymbol{\hat{\theta}}$ and its estimated covariance matrix $\boldsymbol{\Sigma}$ computed with linear regression. Randomization probabilities are then computed using a normal spike-and-slab prior for centered at the origin and with covariance matrix $\boldsymbol{\mathcal{T}}$ with $\boldsymbol{\mathcal{T}}_{ij}=0.5$ for $i\neq j$ and $\boldsymbol{\mathcal{T}}_{ij}=1$ for $i=j$, and for different prior probabilities $\Pr(H_{0})$.
Figure S1 shows sequences of randomization probabilities computed from simulated normal data with $K=3$ treatment groups. We see that assigning a prior probability $\Pr(H_{0})=1$ leads to static equal randomization at $\pi_{i}=1/(K+1)=25\%$, while assigning $\Pr(H_{0})=0$ (Thompson sampling) leads to the most variable randomization probabilities. Since data are simulated assuming that treatment 1 is the most effective, randomization probabilities based on $\Pr(H_{0})<1$ converge towards $\pi_{1}=100\%$ and toward $0\%$ for the remaining treatments. While convergence is the fastest for $\Pr(H_{0})=0$, this prior probability also accidentally produces rather high randomization probabilities for the control group at the start of the study, which is less pronounced for positive prior probabilities $\Pr(H_{0})>0$.
Appendix C Simulation study
@@
C.1 Aims
The aim of the simulation study is to evaluate the design characteristics of the newly proposed null hypothesis Bayesian RAR approach, and compare it to existing methods.
C.2 Data-generating mechanism
@@
Parameters were chosen similar to the simulation study from Robertson et al., (2023). We vary the sample size $n\in\{200,654\}$ to represent low and high powered studies, the number of treatment groups $K\in\{1,2,3\}$, and the probability in the first treatment group $\theta_{1}\in\{0.25,0.35,0.45\}$. The probability in the control group and the remaining groups is always fixed at $\theta_{C}=0.25$ and $\theta_{2}=\theta_{3}=0.3$, respectively. All these parameters are varied fully-factorially, leading to $2\times 3\times 3=18$ parameter conditions.
@@
C.3 Estimands and other targets
@@
The primary interest of this simulation study lies in assessing the patient benefit characteristics of different RAR methods. Additionally, the estimand of interest is the response rate difference $\text{RD}_{1}=\theta_{1}-\theta_{C}$ and the target of interest is the null hypothesis of $\text{RD}_{1}=0$.
C.4 Methods
@@
We consider the null hypothesis Bayesian RAR method described in the main text. The prior probability of $H_{0}$ is a tuning parameter and controls the variability of the randomization probabilities. Setting $\Pr(H_{0})=1$ produces equal randomization, whereas $\Pr(H_{0})=0$ produces Thompson sampling. We consider values of $\Pr(H_{0})\in\{0,0.25,0.5,0.75,1\}$, as well as the normal approximation and exact binomial version of RAR. For approximate normal RAR, a normal prior with mean 0, variance 1, and in case of $K>1$ a correlation of 0.5, is considered. Independent uniform priors are assigned for binomial RAR. Log odds ratios along with their covariance are estimated with logistic regression and then used as inputs for the normal RAR method, while the exact method uses success counts and sample sizes only. In case a method fails to converge, equal randomization is applied as a back-up strategy, as this mimics what an experimenter might do in practice when a RAR method fails to converge (Pawel et al.,, 2025).
@@
We also consider three modifications of these methods: In some conditions, a “burn-in” phase is carried out during which the first 50 patients are always randomized with equal probability $1/(K+1)$ to each group. For Thompson sampling ($\Pr(H_{0})=0$), we additionally consider conditions with power transformations of randomization probabilities, i.e., if $\pi_{k}$ is the randomization probability of group $k$, we take $\pi_{k}^{*}=\pi_{k}^{c}/\sum_{j\in\{C,1,\dots,K\}}\pi_{j}^{c}$. We consider $c=1/2$ and $c=i/(2n)$ with $i$ the current and $n$ the maximal sample size, which are two popular choices of the tuning parameter $c$ (Wathen and Thall,, 2017). Additionally, in some conditions “capping” is applied to Thompson sampling. That is, randomization probabilities outside the $10\%,90\%]$ interval are set to either 10% or 90%. After capping has been performed, randomization probabilities are re-normalized to sum to one (Wathen and Thall,, 2017; Lee and Lee,, 2021). This re-normalization is only performed for randomization probabilities greater than 10%, as these would otherwise be reduced again to probabilities less than 10%. In case, a re-normalized probability becomes less than 10%, it is also capped at 10% and a second re-normalization performed. For equal randomization ($\Pr(H_{0})=1$), no burn-in, capping, or power transformation conditions are simulated as these manipulations have no effect.
Finally, upon suggestions from the associate editor, we also include two methods from the multi-armed bandit literature in the simulation study to compare null hypothesis Bayesian RAR to more aggressive allocation methods – the Gittins index (Gittins et al.,, 2011; Villar et al.,, 2015) and the Bayesian upper confidence bound (Bayes UCB) methods (Kaufmann et al.,, 2012). For the Gittins index, we use the implementation from the RARtrials R package (Xu et al.,, 2025) with a discount factor of $d=0.995$. For Bayes UCB we use uniform priors for each probability along with the tuning parameter $c=0$, as recommended in Kaufmann et al., (2012) based on their simulation results.
C.5 Performance measures
Patient benefit was quantified with:
@@ C.5 Performance measures
    where $n_{1i}$ is the number of allocations to treatment group 1 in simulation $i$. For $K=1$, this reduces to the $\hat{S}_{0.1}$ sample size imbalance measure from Robertson et al., (2023), which in turn was inspired by the performance evaluation in the simulation study from Thall et al., (2015). Note that the $\hat{S}_{0.1}$ measure is difficult to translate into an actual patient benefit measure because the patient benefit loss related to imbalance depends on the size of the treatment effect (Robertson et al.,, 2023). For instance, imbalance is less consequential if the treatment effect is small than if it is large. Nevertheless, we include the $\hat{S}_{0.1}$ measure for comparability with previous simulation studies.
@@
C.6 Computational aspects
The simulation study was run on a server running Debian GNU/Linux forky/sid and R version 4.5.2 (2025-10-31). The SimDesign R package was used to organize and run the simulation study (Chalmers and Adkins,, 2020). Our newly developed brar R package was used to perform Bayesian RAR. Code and data to reproduce this simulation study are available at https://github.com/SamCH93/brar and persistenly archived at https://doi.org/10.5281/zenodo.17248628.
C.7 Results
@@
The mean rate of successes is shown in Figure LABEL:fig:successrate. It was generally the highest for the Gittins index, followed by Bayes UCB, Thompson sampling, and being the lowest for equal randomization. The normal approximation and exact method produced mostly similar rates, with the normal method sometimes showing slightly higher rates (e.g., for $K=3$ and $\text{RD}_{1}=0.2$). The Thompson sampling modifications generally reduced the mean success rate, with the greatest reduction achieved by combining capping and power transformation with $c=i/(2n)$. In conditions with small sample size these rates were similar as when the prior probability was $H_{0}=0.75$, while they were lower when the sample size was larger. This makes sense as for larger sample sizes, uncapped randomization probabilities are more likely to converge to extreme ones. This is also visible in Figure LABEL:fig:extreme where more extreme randomization probabilities are observed with increasing sample size and rate difference for all methods but the ones with capping. Finally, Bayes UCB and the Gittins index methods always produced extreme randomization probabilities, as these methods deterministically allocate the next patient to the group with the highest upper confidence bound for and Gittins index, respectively.
@@
Figure LABEL:fig:imbalance shows negative sample size imbalance as quantified by the $\hat{S}_{0.1}$ metric. The Gittins index generally showed the highest negative imbalance, followed by Thompson sampling and Bayes UCB. Among the methods with $\Pr(H_{0})<1$, negative imbalance was the greatest for Thompson sampling and reduced when modifications were applied. Similarly, increasing the prior probability of $H_{0}$ decreased negative imbalance, in some cases even below Thompson sampling with modifications. As in the simulation study from Robertson et al., (2023), negative imbalance of Thompson sampling and Bayesian RAR with $\Pr(H_{0})<1$ was higher than under equal randomization in conditions with small effect size $\text{RD}_{1}=0.1$, but equal or lower than under equal randomization in conditions with large effect size $\text{RD}_{1}=0.2$.
@@
Figures LABEL:fig:bias and LABEL:fig:coverage show empirical bias and coverage related to estimates of the rate difference between the first treatment group and the control group $\text{RD}_{1}$. In conditions where there was no difference ($\text{RD}_{1}=0$), all methods produced unbiased point estimates, though all methods but equal randomization also showed undercoverage. Bias occurred in conditions with non-zero rate differences, the bias being the greatest for the Gittins index, Bayes UCB, and Thompson sampling, and decreasing to some extent when modifications were introduced or the prior probability of $H_{0}$ increased. Similarly, modifications or increasing prior probabilitities improved coverage, although coverage still remained suboptimal in most conditions. For large rate difference conditions, coverage was often better for the normal than the exact version of RAR. Similarly, bias was in some cases somewhat larger for the exact compared to the normal version. This could potentially be due to the normal prior distribution on the log odds ratio being better tuned to effect sizes in the data-generating mechanism difference compared to the uniform priors on the probabilities under the binomial model.
@@
Figures LABEL:fig:t1e and LABEL:fig:power show empirical type I error rate and power associated with the Wald test of $\text{RD}_{1}=0$. We see that the Gittins index, Bayes UCB, and Thompson sampling methods showed an inflated type I error rate above the nominal 2.5% which was reduced to some extent by increasing either the prior probability of $H_{0}$ or applying modifications. In the same way, Gittins index, Bayes UCB, and Thompson sampling often exhibited reduced power compared to equal randomization, which was again alleviated by modifications or increasing $\Pr(H_{0})$. In small sample sizes and large rate differences, power was slightly increased for the exact compared to the normal version of RAR, while for Thompson sampling the type I rate was slightly higher for the exact compared to the normal version.
Appendix D Software and data
Code and data to reproduce our analyses are openly available at https://github.com/SamCH93/brar. A snapshot of the repository at the time of writing is available at https://doi.org/10.5281/zenodo.17248628. We used the statistical programming language R version 4.5.2 (2025-10-31) for analyses (R Core Team,, 2024) along with the ggplot2 (Wickham,, 2016), dplyr (Wickham et al.,, 2023), SimDesign (Chalmers and Adkins,, 2020), mvtnorm (Genz and Bretz,, 2009), ggh4x (van den Brand,, 2024), ggpubr (Kassambara,, 2023), RARtrials (Xu et al.,, 2025), and knitr (Xie,, 2015) packages.